OpenStudy (.whitedragon.):

Metropolis Gas and Power charges their customers a monthly service charge of $10.50 and a volume charge of $1.64 for every 100 cubic feet of natural gas used. Which of the following equations could be used to find Bruce Wayne's monthly gas bill? (Let x represent the number of cubic feet of gas used in hundreds and y represent the total monthly cost.)

OpenStudy (.whitedragon.):

y = 8.86x y = 12.14x y = 1.64x + 10.50 y = 1.64x

OpenStudy (anonymous):

lol bruce wayne

OpenStudy (.whitedragon.):

...-_-...

OpenStudy (anonymous):

its c

OpenStudy (.whitedragon.):

The table above represents what type of function? neither linear nor nonlinear linear and nonlinear nonlinear linear

OpenStudy (.whitedragon.):

x y 0 -1 1 1 2 3 3 5

OpenStudy (anonymous):

whats linear function and non linear function

OpenStudy (.whitedragon.):

A linear problem is any problem which is solved by setting up only linear equations or linear systems of equations to solve. An expression in the variables x1,...,xn is linear if it is of the form a1x1+...+anxn+b. Here a1,...an and b are coefficients. If the expressions on both sides of the equation are of this form then it is a linear equation. If a problem requires nonlinear equations to solve then it is a nonlinear problem. For example xy=1 is not linear, as is e^x+sin(y)=ln(x+y) and may only arrise in solving nonlinear problems.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

plot the points c if they give you a straight line if they do then its a linear function